# LouisMahon

### v1

@@ 1 Introduction
There exist several classic approaches to measuring complexity. Kolmogorov complexity takes the length of the shortest program that generates the given data. It is uncomputable, but there are computable approximations, such as file-compression ratio under some compression algorithm. A similar approach is found in the minimum description length principle [1] (MDL), though this is mostly concerned with fitting models. Entropy is closely related to MDL, and is often used as a measure of complexity for data of various sorts [2, 3]. One problem that arises with these methods, and their many derivatives, is that they give a low score to simple, highly regular data, and a high score to random, noisy data, with the data we consider most meaningful, such as human speech, falling somewhere between the two. Thus, we can say neither that a high score nor a low score indicates the data is meaningful. This fundamental issue has been identified by various authors [4, 5, 6, 7, 8].
The basis for our method is that length of the smallest representation of a piece of data indicates how complex it is. However, unlike prior methods, either algorithmic such as Kolmogorov complexity, or statistical such as entropy, we make a division of the description into a ‘meaningful’ and ‘meaningless’ portion. This is similar, on a high level, to some theoretical work to divide the Kolmogorov complexity into meaningful information and noise, using, e.g., ‘sophistication’ [8, 6] or ‘effective complexity’ [5, 7]. When selecting the shortest overall description of the data, we include both the meaningful and the meaningless portion, but, after making this selection, we ignore the meaningless portion and take only the length of the meaningful portion as contributing to the complexity. Put semi-formally, if $d$ is a description of data $X$, and $m(d)$ is the meaningful portion of $d$, then our proposed meaningfulness score is given by $|m(d^{*})|$, where $d^{*}=\operatorname*{argmin\,}_{d}|d(X)|$. Note that the meaningless portion still plays an important role, as it helps select the optimal description.

### v2

@@ 1 Introduction
There exist several classic approaches to measuring complexity. Kolmogorov complexity (KC) takes the length of the shortest program that generates the given data. It is uncomputable, but can be approximated by file-compression ratio. The minimum description length principle [1] (MDL) principle is an information-theoretic approach to the same idea that the one should search for the shortest description of the data. Mostly, MDL is concerned with model selection, though has also been used for other applications such as a parameter-free algorithm for optimal segmentation [2]. [3] has used MDL for measuring complexity. Entropy is closely related to MDL, and is often used as a measure of complexity for data of various sorts [4, 5]. One problem that arises with these methods, and their many derivatives, is that they give a low score to simple, highly regular data, and a high score to random, noisy data, with the data we consider most meaningful, such as human speech, falling somewhere between the two. Thus, we can say neither that a high score nor a low score indicates the data is meaningful. This fundamental issue has been identified by various authors [3, 6, 7, 8, 9].
We follow the MDL principle for our method, but make a novel division of the description into a ‘meaningful’ and ‘meaningless’ portion. This is similar, on a high level, to some theoretical work to divide KC into meaningful information and noise, using, e.g., ‘sophistication’ [9, 7] or ‘effective complexity’ [6, 8]. When selecting the shortest description of the data, we include both the meaningful and the meaningless portion, but, after making this selection, we ignore the meaningless portion and take only the length of the meaningful portion as contributing to the complexity. Put semi-formally, if $d$ is a description of data $X$, and $m(d)$ is the meaningful portion of $d$, then our proposed meaningfulness score is given by $|m(d^{*})|$, where $d^{*}=\operatorname*{argmin\,}_{d}|d(X)|$. Note that the meaningless portion still plays an important role, in selecting the optimal description.